# Environmentally Sustainable Software Design and Development: A Systematic Literature Review

### v1

@@ Abstract.
The ICT sector, responsible for 2% of global carbon emissions and significant energy consumption, is under scrutiny calling for methodologies and tools to design and develop software in an environmentally sustainable-by-design manner. However, the software engineering solutions for designing and developing sustainable software are currently scattered over multiple different pieces of literature, which makes it difficult to consult the body of knowledge on the topic. In this article, we precisely conduct a systematic literature review on state-of-the-art proposals for designing and developing sustainable software. We identify and analyse 65 primary studies by classifying them through a taxonomy aimed at answering the 5W1H questions of environmentally sustainable software design and development. We first provide a reasoned overview and discussion of the existing guidelines, reference models, measurement solutions and techniques for measuring, reducing, or minimising the energy consumption and carbon footprint of software. Ultimately, we identify open challenges and research gaps, offering insights for future work in this field.
@@ 1. Introduction
In recent years, the themes of sustainability and sustainable development have seen a growing interest from governments, academia and industry, aiming at balancing the need to safeguard our planet along with social and economical instances (Brundtland et al., 1987; Sachs et al., 2022). Such effort is clearly epitomised by the adoption of the Sustainable Development Goals in the UN action plan reported by the 2030 Agenda for Sustainable Development.111Publicly available at: https://sdgs.un.org/2030agenda Indeed, global carbon emissions need to be reduced to secure our Planet’s future by achieving the ambition set by the Paris agreement.222https://www.un.org/en/climatechange/paris-agreement However, such an objective remains far from being reached, as highlighted by the outcomes of the 2023 United Nations Climate Change Conference (COP28).
@@
Building on this momentum, there is a rising concern about measuring, assessing, and enhancing the sustainability of Information and Communication Technology (ICT). Indeed, the ICT is currently responsible for 2% of the global carbon emissions, i.e. as much as airplane traffic, with an energy consumption estimated at between 6% and 9% of the global demand and expected to grow up to 20% by 2030 (Belkhir and Elmeligi, 2018; Zulfiqar et al., 2023; Adisa et al., 2024). Most of the efforts to reduce the carbon footprint of ICT were dedicated to improve hardware energy-efficiency, by leveraging the combined effects of Moore’s law (Moore, 1998) and Dennard scaling (Dennard et al., 1974). Moore’s law ensured the possibility of packing double the transistors in a chip approximately every 18 months, while Dennard scaling has ensured that integrated circuits of the same area maintain constant power consumption regardless of the increased number of transistors. This resulted in faster chips without an increase in energy consumption up until the 2010s (Anderson et al., 2023). Thus, Cloud providers would upgrade their infrastructure biennially to double computing power without increasing energy costs and reducing their carbon emissions per unit of computation.
@@
With Dennard’s scaling and Moore’s law reaching their limits (Muralidhar et al., 2022), improving the sustainability of ICT will mandatorily involve making software more sustainable by reducing (i) the energy consumption and/or (ii) the carbon footprint associated with its deployment. On this line, much research has targeted the optimisation of renewable energy usage through suitable service orchestration and consolidation (Gaglianese et al., 2023a), without considering the potential of (re)designing and (re)writing software following a sustainable-by-design approach. Naturally, as pointed out by Anderson et al. (Anderson et al., 2023), further improvements towards improving software sustainability will require treating “energy and carbon as a first-class resource” in our code design and development. Similarly, Verdecchia et al. (Verdecchia et al., 2021) call for methodologies to write sustainable software by blending decision-making with the software execution context. Although the term green software engineering was coined more than a decade ago (Caballar, 2024), much remains to be done to fully rethink how software is designed and implemented in an environmentally sustainable manner.
@@
This article precisely aims at surveying the state-of-the-art related to the design and development of environmentally sustainable software. The currently available sustainable software engineering methodologies are indeed scattered over multiple different pieces of literature. Unfortunately, this makes it difficult to consult the body of knowledge on the topic, both for researchers willing to investigate on sustainable software engineering methodologies and for practitioners willing to apply them to their software systems. To achieve this objective, this article conducts a systematic literature review (SLR), which identifies, taxonomically classifies, and compares the existing guidelines, reference models, measurement solutions, and techniques for designing and developing sustainable software. Following the guidelines by Petersen et al. (Petersen et al., 2008; Petersen et al., 2015) on how to conduct SLRs in software engineering, we first identify a corpus of 65 primary studies, dating from 2001 onwards. We then classify and compare the selected studies by means of a newly proposed taxonomy, aimed at highlighting the 5W1H of sustainable software engineering, viz., who are the stakeholders targeted by the selected studies, what methods are proposed therein, why, where and when the proposed methods are applied (in terms of application domain and software lifecycle step, respectively), and how the selected studies have been released to the community.
@@
Overall, our SLR pursues a twofold ultimate objective. On one hand, it aims at accurately depicting existing proposals, methodologies, and solutions that allow the design and development of sustainable software, namely for measuring, reducing, and/or minimising a software’s energy consumption and carbon emissions. On the other hand, it analyses the state of the art on sustainable software engineering with the aim of shedding light on open challenges and gaps, which can contribute to reducing the carbon footprint of ICT from the software perspective.
@@
The rest of this article is organised as follows. Section 2 describes the design of the SLR process. Section 3 classifies and recaps the state of the art on environmentally sustainable software design and development. The classification is retaken in Section 4, which provides an analysis of the 5W1H of designing and developing environmentally sustainable software. Sections 5 and 6 discuss the possible threats to the validity of our SLR and related work, respectively. Finally, Section 7 concludes the article by shedding light on future research directions for the design and development of environmentally sustainable software.
@@ 2. Research Design
We also provide a replication package for repeating our SLR (Section 2.5)
@@ 2.1. The 5W1H Questions
The objective of this article is to identify, classify, and analyse the state of the art on the design and development of environmentally sustainable software by limiting its energy consumption and/or carbon footprint. This is done by answering to the 5W1H questions listed below:
Who? Who are the stakeholders for the existing techniques for sustainable software design and development?
What? What methods, tools, and techniques have been proposed to achieve sustainable software design and development?
Why? Why is a selected study targeting the environmental sustainability of software? Is it aimed at reducing its energy consumption, carbon footprint, or both?
@@ 2.2. Literature Search
Following the guidelines by Petersen et al. (Petersen et al., 2008; Petersen et al., 2015), we identified the search string guided by our main research objective, namely to overview the state of the art on sustainable software design and development. More precisely, we defined the search strings based on the PICO (Population, Intervention, Comparison, Outcome) terms characterising our research questions (Kitchenham and Charters, 2007), taking keywords from each aspect of our research objective – and its corresponding 5W1H questions. Differently from Petersen et al. (Petersen et al., 2008; Petersen et al., 2015), we did not restrict our focus to specific outcomes or experimental designs in our study. We indeed aimed at providing a broader overview of the state-of-the-art for the design and development of environmentally sustainable software. This difference is also reflected by the search strings we employed to search for studies, namely,
@@
We used the aforementioned string to search for white literature published until December 2023. The search was carried on ACM Digital Library, IEEEXplore, Scopus, and Web of Science. We retrieved 1416 candidate studies overall, which – after checking for duplicates – were reduced to a list of 1250 unique candidate studies.
@@ 2.3. Literature Selection
The search terms used in Section 2.2 were checked against the title and abstract of existing literature, obtaining a significant number of candidate studies that were not relevant to our purposes. Therefore, the list of candidate studies was further refined by relying on the inclusion/exclusion criteria in Table 1. The inclusion criteria and aligned the selection with our main research objective, by allowing to select the existing studies that support the design and/or the development of software, while also aiming at reducing its carbon footprint. The exclusion criteria , , and restricted the focus to peer-reviewed literature written in English, which was also accessible from our institution. Finally, the exclusion criteria and restricted the focus to primary studies, by excluding collection of papers and secondary studies, respectively. A candidate study was selected if satisfying all the inclusion criteria, whereas it was excluded if satisfying any of the exclusion criteria.
Criteria | Type | Description |
@@
Inclusion | The study proposes solutions to support the design and/or development of software | |
@@
Exclusion | The study is a collection of papers (e.g., conference proceedings) | |
@@
The inclusion/exclusion criteria were applied in two subsequent steps. In both steps, each candidate study was marked as to be selected or excluded by two authors, independently. In case of disagreement on the inclusion of a candidate, the author who did not participated in the coding of the study acted as a third evaluator to resolve the disagreement. In the first step, the selection was based on assessing the inclusion/exclusion criteria over the title and abstract of candidate studies, and this reduced the list of candidate studies to 88 candidate studies, with an overall inter-rater agreement of 84.4%.333We measured the inter-rater agreement on the inclusion/exclusion coding by exploiting the Krippendorff- coefficient (Krippendorff, 2004), which classifies a coding as good if the value is greater or equal to 80%. Then, the process was repeated by considering the full text of the remaining 88 candidate studies, and this further narrowed the list of candidate studies to 43 candidate studies, with an overall inter-rater agreement of 90.9%.
Ref. Author(s). Title (Year) |
@@
(Alofi et al., 2023) Alofi et al. Self-optimizing the environmental sustainability of blockchain-based systems (2023) |
@@
(Atkinson et al., 2014) Atkinson et al. Facilitating greener IT through green specifications (2014) |
@@
(Beghoura et al., 2014) Beghoura et al. Green application awareness: nonlinear energy consumption model for green evaluation (2014) |
@@
(Chauhan and Saxena, 2013) Chauan and Saxena. A green software development life cycle for cloud computing (2013) |
@@
(Dick et al., 2013) Dick et al. Green software engineering with agile methods (2013) |
@@
(Fakhar et al., 2012) Fakhar et al. Software level green computing for large scale systems (2012) |
@@
(Goaër, 2021) Le Goaër. Enforcing green code with Android lint (2021) |
@@
(Goaër and Hertout, 2022) Le Goaër. ecoCode: a SonarQube plugin to remove energy smells from Android projects (2023) |
@@
(Hoesch-Klohe et al., 2010) Hoesch-Klohe et al. Towards green business process management (2010) |
@@
(Ibrahim et al., 2022) Ibrahim et al. Green software process factors: a qualitative study (2022) |
@@
(Jebraeil et al., 2017) Jebraeil et al. gUML: Reasoning about energy at design time by extending UML deployment diagrams with data centre contextual … (2017) |
@@
(Karunakaran and Rao, 2013) Karunakaran and Rao. A petri net simulation of software development lifecycle towards green IT (2013) |
@@
(Kern et al., 2015) Kern et al. Impacts of software and its engineering on the carbon footprint of ICT (2015) |
@@
(Kharchenko and Gorbenko, 2013) Kharchenko and Gorbenko. Evolution of von Neumann’s paradigm: dependable and green computing (2013) |
@@
(Kipp et al., 2012) Kipp et al. Layered green performance indicators (2012) |
@@
(Lago, 2019) Lago. Architecture design decision maps for software sustainability (2019) |
@@
(Lami and Buglione, 2012) Lami and Buglione. Measuring software sustainability from a process-centric perspective (2012) |
@@
(Lami et al., 2013) Lami et al. A methodology to derive sustainability indicators for software development projects (2013) |
@@
(Mahmoud and Ahmad, 2013) Mahmoud and Ahmad. A green model for sustainable software engineering (2013) |
@@
(Mehra et al., 2023) Mehra et al. Assessing the impact of refactoring energy-inefficient code patterns on software sustainability: an industry case study (2023) |
@@
(Mohankumar and Kumar, 2016) Mohankumar and Kumar. Green based software development life cycle model for software engineering (2016) |
@@
(Moshnyaga, 2013b) Moshnyaga. Assessment of software lifecycle energy and its contribution to green house gas emissions (2013) |
@@
(Naumann et al., 2011) Naumann et al. The GREENSOFT model: a reference model for green and sustainable software and its engineering (2011) |
@@
(Noureddine et al., 2012) Noureddine et al. A preliminary study of the impact of software engineering on GreenIT (2012) |
@@
(Pan and Rajan, 2022) Pan and Rajan. Decomposing convolutional neural networks into reusable and replaceable modules (2022) |
@@
(Penzenstadler and Femmer, 2013) Penzenstadler and Femmer. A generic model for sustainability with process– and product-specific instances (2013) |
@@
(Radersma, 2022) Radersma. Green coding: reduce your carbon footprint (2022) |
@@
(Sharma et al., 2015) Sharma et al. Energy efficient software development life cycle – an approach towards smart computing (2015) |
@@
(Siegmund et al., 2022) Siegmund et al. Green configuration: can artificial intelligence help reduce energy consumption of configurable software systems? (2022) |
@@
(Simon et al., 2023) Simon et al. Uncovering the environmental impact of software life cycle (2023) |
@@
(Tee et al., 2014b) Tee et al. Web-based knowledge management model for managing and sharing green knowledge of software development in … (2014) |
@@
(Wei et al., 2023) Wei et al. Towards greener yet powerful code generation via quantization: an empirical study (2023) |
@@
The list of candidate studies was completed by running backward snowballing (Wohlin, 2014), which resulted in 30 additional candidate studies. These were again classified for inclusion/exclusion by repeating the same process as described above – and by directly considering their full read. As a result, 22 additional studies were selected, reaching a total of 65 studies to be considered by our SLR. The final list of selected studies is in Table 2.
@@ 2.4. Literature Classification
The selected studies were classified by exploiting the taxonomy in Table 3. The taxonomy is rooted in the 5W1H questions driving our research (Section 2.1), with each of such questions being mapped to one or more possible classifications, i.e., categories of values. For the sake of conciseness, for each category, Table 3 only lists the values that were associated to at least one selected study.
5W1H | Category | Values | |
@@
What? | proposal | sustainability guidelines | experience report, good practices |
reference models | reference model | ||
measurement solutions | assistive tooling, estimate technique | ||
software improvement techniques | decision making, distributed protocol, ML, model-driven, optimisation, search-based software engineering, static analysis, testing | ||
source code availability | black-box, white-box | ||
TRL | 1–9 | ||
Why? | goal* | energy | measure, reduce, optimise |
carbon emissions | measure, reduce, optimise | ||
When? | lifecycle stage* | requirement analysis, design, implementation, integration & testing, rollout | |
@@
How? | publication year | 2001–2023 | |
publication type | book chapter, conference paper, journal article |
@@
Who? The taxonomy enables classifying the (potentially multiple) stakeholders for the selected studies. These include the end-user (who selects and uses a software product), the IT operator (who operates a software product), the product owner (who handles the software process and aims at maximising the value of a software product), the software architect (who designs of a software product), the software developer (who implements parts of the software product), the software provider (who delivers a software product to end-users), and the sustainability engineer (who engineers the sustainability of a software product).
@@
What? The taxonomy in Table 3 enables classifying the proposal of each selected study, by distinguishing sustainability guidelines (in the form of shared good practices or experience reports), reference models (namely, abstract frameworks consisting of interlinked concepts aimed at clarifying what is sustainaibility for software design and development), measurement solutions (in the form of assistive tooling or estimate techniques supporting the measurement of a software’s sustainability), and software improvement techniques (namely, techniques aimed at supporting the design and/or development of software to enhance its environmental sustainability).
@@
The taxonomy also enables distinguishing whether a selected study’s proposal is white-box (or black-box), namely whether it requires the availabilty of a software’s source code (or not). It is also possible to indicate the TRL (Technology Readiness Level) of a selected study’s proposal from 1 to 9, following the TRL classification defined by the European Commission (European Commission, 2014).
@@
Why? The taxonomy in Table 3 enables classifying the (potentially multiple) goals of the selected studies. Namely, it enables indicating whether a selected study aims at supporting the design and development of software while measuring, reducing, and/or optimising its energy consumption/carbon emissions. While the reduction and optimisation of energy consumption/carbon emission readily enhance the environmental sustainability of a software, measurement can also support it, e.g., by allowing to measure a software’s environmental sustainability before and after updates, to check whether such updates enhanced/deteriorated the software’s sustainability.
@@
When? The taxonomy in Table 3 enables distinguishing the (potentially multiple) lifecycle stages supported by a selected study’s proposal. To this end, we follow the iterative software lifecycle process described by Kneuper (Kneuper, 2018), which consists of requirement analysis, design, implementation, integration & testing, and rollout, in this order.
@@
Where? The taxonomy in Table 3 enables distinguishing the target application domain for a selected study. These include blockchains, cloud computing, computer vision, cyber-physical systems, edge computing, HPC, ML, and mobile applications. The above listed domains are complemented by the generic software value, which is used when the proposal is intended to work with any software – rather than targeting and exploting the peculiarities of a specific application domain.

### v2

# Carbon-Efficient Software Design and Development: A Systematic Literature Review

@@ Abstract.
The ICT sector, responsible for 2% of global carbon emissions, is under scrutiny calling for methodologies and tools to design and develop software in an environmentally sustainable-by-design manner. However, the software engineering solutions for designing and developing carbon-efficient software are currently scattered over multiple different pieces of literature, which makes it difficult to consult the body of knowledge on the topic. In this article, we precisely conduct a systematic literature review on state-of-the-art proposals for designing and developing carbon-efficient software. We identify and analyse 65 primary studies by classifying them through a taxonomy aimed at answering the 5W1H questions of carbon-efficient software design and development. We first provide a reasoned overview and discussion of the existing guidelines, reference models, measurement solutions and techniques for measuring, reducing, or minimising the carbon footprint of software. Ultimately, we identify open challenges and research gaps, offering insights for future work in this field.
@@ 1. Introduction
In recent years, the themes of sustainability and sustainable development have seen a growing interest from governments, academia and industry, aiming at balancing the need to safeguard our planet along with social and economical instances (Brundtland et al., 1987; Sachs et al., 2022). Such effort is clearly epitomised by the adoption of the Sustainable Development Goals in the UN action plan reported in the 2030 Agenda for Sustainable Development.111Publicly available at: https://sdgs.un.org/2030agenda Indeed, global carbon emissions need to be reduced to secure our Planet’s future by achieving the ambition set by the Paris agreement222https://www.un.org/en/climatechange/paris-agreement (Manner, 2022; Möller et al., 2024). However, such an objective remains far from being reached, as highlighted by the outcomes of the 2023 United Nations Climate Change Conference (COP28).
@@
Building on this momentum, there is a rising concern about measuring, assessing, and improving the sustainability of Information and Communication Technology (ICT). Indeed, ICT is currently responsible for 2% of global carbon emissions (Manner, 2022; Mingay and Pamlin, 2008), with an energy consumption estimated at between 6% and 9% of the global demand and expected to grow up to 20% by 2030 (Belkhir and Elmeligi, 2018; Zulfiqar et al., 2023; Adisa et al., 2024). Most efforts to reduce the carbon footprint of ICT were dedicated to improve hardware energy-efficiency, by leveraging the combined effects of Moore’s law (Moore, 1998) and Dennard scaling (Dennard et al., 1974). Moore’s law ensured the possibility of packing double the transistors in a chip, while Dennard scaling has ensured that integrated circuits of the same area maintain constant power consumption regardless of the increased number of transistors. This resulted in faster chips with no increase in energy consumption up until the 2010s (Anderson et al., 2023).
@@
With Dennard’s scaling and Moore’s law reaching their limits (Muralidhar et al., 2022), improving the environmental sustainability of ICT will necessarily involve making software more sustainable by reducing the carbon footprint associated with its deployment. On this line, much research has targeted the optimisation of renewable energy usage through suitable service orchestration and consolidation (Gaglianese et al., 2023a), without considering the potential of (re)designing and (re)writing software according to a sustainable-by-design approach. Naturally, as pointed out by Anderson et al. (Anderson et al., 2023), further improvements in software sustainability will require treating carbon-efficiency in our code design and development. Similarly, Verdecchia et al. (Verdecchia et al., 2021) call for methodologies to write sustainable software by blending decision-making with the software execution context. Although the term green software engineering was coined more than a decade ago (Caballar, 2024), much remains to be done to fully rethink how carbon-efficient software is designed and implemented.
At the same time, as observed, e.g., by the Green Software Foundation (Murugesan, 2021) and Kern et al. (Kern et al., 2015), the carbon-efficiency of software products is mostly affected by their design, development, and operation. Indeed, while sustainability considerations span the entire software lifecycle, design, development, and operation represent the primary leverage points for reducing software carbon emissions, e.g., through software architecture decisions, efficient data structures and algorithms, and energy- and carbon-aware orchestration strategies.
@@
While the environmentally sustainable operation of software products has already been analysed, e.g., by Gaglianese et al. (Gaglianese et al., 2023b), this article precisely aims at surveying the state of the art related to the design and development of carbon-efficient software. Currently available methodologies are scattered over multiple different pieces of literature. Unfortunately, this makes it difficult to consult the body of knowledge on the topic, both for researchers willing to investigate carbon-efficient software design and development methodologies, and for practitioners willing to apply them to their software systems. To achieve this objective, this article conducts a systematic literature review (SLR), which identifies, taxonomically classifies, and compares the existing guidelines, reference models, measurement solutions, and techniques for designing and developing carbon-efficient software. Following the guidelines by Petersen et al. (Petersen et al., 2008, 2015) on how to conduct SLRs in software engineering, we first identify a corpus of 65 primary studies. We then classify and compare the selected studies by means of a newly proposed taxonomy, aimed at highlighting the 5W1H of carbon-efficient software design and development, viz., who are the stakeholders targeted by the selected studies, what methods are proposed therein, why, where and when the proposed methods are applied (in terms of application domain and software lifecycle step, respectively), and how the selected studies have been released to the community.
@@
Overall, our SLR pursues a twofold ultimate objective. On one hand, it aims at accurately depicting existing proposals, methodologies, and solutions that enable the design and development of environmentally sustainable software, namely for measuring, reducing, and/or minimising a software’s carbon emissions. On the other hand, it analyses the state of the art in carbon-efficient software design and development with the aim of shedding light on open challenges and gaps, which can contribute to reducing the carbon footprint of ICT from a software perspective.
@@
The rest of this article is organised as follows. Section 2 describes the design of the SLR process. Section 3 classifies and summarises the state of the art on carbon-efficient software design and development. The classification is retaken in Section 4, which provides an analysis of the 5W1H of designing and developing carbon-efficient software. Sections 5 and 6 discuss the possible threats to the validity of our SLR and related work, respectively. Finally, Section 7 concludes the article by highlighting future research directions for the design and development of carbon-efficient software.
@@ 2. Research Design
We also provide a replication package for repeating our SLR (Section 2.5).
@@ 2.1. The 5W1H Questions
The objective of this article is to identify, classify, and analyse the state of the art in the design and development of environmentally sustainable software by limiting its carbon footprint. This is done by answering to the 5W1H questions listed below:
Who? Who are the stakeholders in existing techniques for carbon-efficient software design and development?
What? What methods, tools, and techniques have been proposed to achieve carbon-efficient software design and development?
Why? Why does a selected study target the carbon efficiency of software? Does it aim at measuring, reducing, or optimising its carbon footprint?
@@ 2.2. Literature Search
Following the guidelines by Petersen et al. (Petersen et al., 2008, 2015), we identified the search string guided by our main research objective, namely to overview the state of the art in carbon-efficient software design and development. More precisely, we defined the search strings based on the PICO (Population, Intervention, Comparison, Outcome) terms characterising our research questions (Kitchenham and Charters, 2007), taking keywords from each aspect of our research objective – and the corresponding 5W1H questions. Differently from Petersen et al. (Petersen et al., 2008, 2015), we did not limit our focus to specific outcomes or experimental designs in our study. Rather, we aimed at providing a broader overview of the state of the art in the design and development of carbon-efficient software. This difference is also reflected in the search string we employed to search for studies, i.e.,
@@
We used the aforementioned string to search for all white literature published before January 2024, when this research began. The search was carried on ACM Digital Library, IEEEXplore, Scopus, and Web of Science, and was restricted to white literature published in Computer Science venues. We retrieved a total of 1517 candidate studies, which – after checking for duplicates – were reduced to a list of 1273 unique candidate studies.
@@ 2.3. Literature Selection
The search terms used in Section 2.2 were checked against the titles and abstracts of the existing literature, which resulted in a significant number of candidate studies that were not relevant for our purposes. Therefore, the list of candidate studies was further refined by relying on the inclusion/exclusion criteria in Table 1. The inclusion criteria and aligned the selection with our main research objective by allowing the selection of existing studies that support the design and/or development of software, while also aiming at reducing its carbon footprint. The exclusion criteria , , and restricted the focus to peer-reviewed literature written in English, which was also accessible from our institution. Finally, the exclusion criteria and restricted the focus to primary studies and excluding collections of papers and secondary studies, respectively. A candidate study was selected if satisfying all the inclusion criteria, whereas it was excluded if satisfying any of the exclusion criteria.
[…]
@@
[…]
@@
[…]
@@
The inclusion/exclusion criteria were applied in two consecutive steps. In both steps, each candidate study was marked as to be selected or excluded by two authors independently. If there was disagreement about the inclusion of a candidate study, the author who was not involved in the coding of the study acted as a third evaluator to resolve the disagreement. In the first step, the selection was based on assessing the inclusion/exclusion criteria over the title and abstract of candidate studies, which reduced the list of candidate studies to 88 candidate studies, with an overall inter-rater agreement of 85.6%.333We measured the inter-rater agreement on the inclusion/exclusion coding by exploiting the Krippendorff- coefficient (Krippendorff, 2004), which classifies a coding as good if the value is greater or equal to 80%. The process was then repeated by considering the full text of the remaining 88 candidate studies, which further narrowed the list to 43 candidate studies, with an overall inter-rater agreement of 90.9%.
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The list of candidate studies was completed by running backward snowballing (Wohlin, 2014), which resulted in 30 additional candidate studies. These were again classified for inclusion/exclusion by repeating the same process as described above and by directly considering their full read. As a result, 22 additional studies were selected, reaching a total of 65 studies to be considered by our SLR. The final list of selected studies is shown in Table 2.
@@ 2.4. Literature Classification
The selected studies were classified by exploiting the taxonomy in Table 3. The taxonomy is rooted in the 5W1H questions driving our research (Section 2.1), with each of these questions being mapped to one or more possible classifications, i.e., categories of values. For the sake of brevity, Table 3 lists for each category the values associated with at least one selected study.
[…]
@@
[…]
| Why? | goal* | indirectly | measure, reduce, optimise |
| directly | measure, reduce, optimise | ||
[…]
@@
[…]
@@
Who? The taxonomy enables classifying the (potentially multiple) stakeholders for the selected studies. These include the end-user (who selects and uses a software product), the IT operator (who operates a software product), the product owner (who manages the software process and aims at maximising the value of a software product), the software architect (who designs of a software product), the software developer (who implements parts of a software product), the software provider (who delivers a software product to end-users), and the sustainability engineer (who engineers the sustainability of a software product).
@@
What? The taxonomy in Table 3 enables classifying the proposal of each selected study by distinguishing among sustainability guidelines (in the form of shared good practices or experience reports), reference models (namely, abstract frameworks consisting of interlinked concepts aimed at clarifying what is carbon-efficiency for software design and development), measurement solutions (in the form of assistive tooling or estimate techniques supporting the measurement of a software’s sustainability), and software improvement techniques (namely, techniques aimed at supporting the design and/or development of software to reduce its carbon footprint).
@@
The taxonomy also enables distinguishing whether the proposal of a selected study is white-box (or black-box), namely whether it requires the availabilty of a software’s source code, given its relevance to software design and development. It is also possible to indicate the TRL (Technology Readiness Level) of a selected study’s proposal from 1 to 9, following the TRL classification defined by the European Commission (European Commission, 2014).
@@
Why? The taxonomy in Table 3 enables classifying the (potentially multiple) goals of the selected studies. Namely, it enables indicating whether a selected study aims at supporting the design and development of software while measuring, reducing, or optimising its carbon emissions. While reducing and optimising carbon emissions will certainly improve a software’s carbon efficiency, measurement can also support this, for example by comparing a software’s carbon footprint before and after updates. This is to check whether such updates have made the software more or less sustainable. We also distinguish whether this is done by directly measuring a software’s carbon emissions, or indirectly, by assuming that reducing a software’s energy consumption will also reduce its carbon footprint.
@@
When? The taxonomy in Table 3 enables distinguishing the (potentially multiple) lifecycle stages supported by a selected study’s proposal. To this end, we follow the iterative software lifecycle process described by Kneuper (Kneuper, 2018), which consists of requirement analysis, design, implementation, integration & testing, and rollout, in that order.
@@
Where? The taxonomy in Table 3 makes it possible to distinguish the target application domain for a selected study. These include blockchains, cloud computing, computer vision, cyber-physical systems, edge computing, HPC, ML, and mobile applications. The above listed domains are complemented by the generic software value, which is used when the proposal is intended to work with any software, rather than targeting and exploiting the peculiarities of a specific application domain.